# Deep Metric Imitation Learning for Stable Motion Primitives

### v2

@@ 5.2.1 Accuracy
Fig. 10a presents the performance of two variations of PUMA, namely with Euclidean and spherical distance functions. Moreover, BC is included, serving as an upper-bound reference. As explained in Sec. 5.1.1, the boundary loss $\ell_{\partial}$ does not apply in this case; hence, it is not evaluated. We can observe that the accuracy performance of both variations of PUMA is very similar, and BC performs slightly better than both of them.
@@ 5.2.2 Stability
Table 3 depicts the stability results for the LASA $\mathcal{S}^{2}$ dataset. In this case, we also conducted a stability analysis of CONDOR, as our discussion in Sec. LABEL:sec:non-euclidean suggested that this method should not be capable of ensuring stability in non-Euclidean state spaces. This assertion is confirmed by the data in Tab. 3, which shows that CONDOR yields 7.3133% of unsuccessful trajectories. When compared to BC, which has a 15.9217% rate of unsuccessful trajectories, it can be inferred that CONDOR reduces the number of unsuccessful trajectories in non-Euclidean state spaces. However, its performance is still far from satisfactory. In contrast, both versions of PUMA achieve 0% of unsuccessful trajectories, marking a significant improvement. Moreover, it validates that the Euclidean distance is effective for enforcing stability in spheres, as it can induce spherical metrics, such as the chordal distance, in lower-dimensional manifolds.
@@ 5.2.3 Qualitative Analysis
Fig. 11 illustrates the motions learned with PUMA in $\mathcal{S}^{2}$, using both spherical and Euclidean distance. Within the observed region of the sphere, the vector field exhibits no spurious attractors. Furthermore, when initiated from the same conditions as the demonstrations, the trajectories in both cases show an accurate reproduction of the motions.
@@ 5.2.4 State-of-the-art Comparison
Fig. 10b depicts a comparison between three methods: 1) CONDOR, 2) PUMA, and 3) a method developed for learning stable motions in non-Euclidean manifolds using tools from Lie Theory, which we will refer to as Lie Flows [11]. PUMA outperforms both CONDOR and Lie Flows in terms of accuracy. Moreover, in contrast to CONDOR, PUMA also successfully ensured stability in this dataset. This clearly shows PUMA’s superiority over CONDOR in non-Euclidean state spaces.
@@ 5.3 Real-world Experiments
We validate our method in two real-world setups, using two different robots. Both robots have six degrees of freedom, with their end-effector pose represented in $\mathbb{R}^{3}\times\mathcal{S}^{3}$ and controlled using PUMA. In these experiments, $f_{\theta}^{\mathcal{T}}$ is employed to generate velocity references in real time, which are then sent to the low-level controllers responsible for tracking these references in the robots. Note that throughout the experiments, we operated under the assumption that the target states of the robots are already known.
@@ 5.3.1 Greenhouse
This experiment was conducted in a greenhouse, where a robot was trained to reach a black marker on a tomato plant (see Fig. 12). This marker, simulating a plant’s peduncle, represented the target state the robot must reach to harvest the plant. Utilizing the marker in this experiment allowed us to test the method without causing harm to the plant. This task presents significant challenges, as it requires the robot to perform precise position and orientation control while considering the plant’s complex geometry, which is non-trivial to model. In a practical setting, a motion library should be developed to account for the unique characteristics of individual plants and targets. This library could then be used to select or combine suitable motions to create an appropriate movement strategy that considers these factors. This experiment represents a preliminary step towards achieving this objective.
In this experiment, we utilized an ABB IRB 1200 robot, which can be equipped with the Externally Guided Motion [^7] Code: https://github.com/ros-industrial/abb_robot_driver. module, allowing real-time joint velocity control. As the motion takes place in the end-effector space, the desired velocity at each time step was calculated using $f_{\theta}^{\mathcal{T}}$, and subsequently mapped to the joint space. This conversion was accomplished using the inverse kinematics module with joint limits from the Robotics Toolbox [50]. Similarly, we translated the robot’s estimated joint state into the end-effector space using the forward kinematics module from the same toolbox. The same approach was taken to collect the demonstrations of this task directly in the end-effector space, where a space mouse was employed to teleoperate the robot.
Fig. 13 presents the demonstrations and simulations of motions conducted in this experiment. It can be observed that five demonstrations were provided. Note that these demonstrations do not need to have the same end time; only the final state needs to be consistent. Over a span of 70 seconds, all simulated trajectories visibly converge to the goal. For more details, the reader is referred to the attached video.
@@ 5.3.2 Hammer
The second experiment involves a robot accurately positioning a hammer next to other tools on a table. This task requires precise control of the robot’s position and orientation, as shown in Fig. 12. We are interested in controlling the hammer’s movement, so we assume that the state of the hammer directly relates to the state of the robot’s end effector. This assumption is necessary because the employed low-level controller acts on the robot’s end-effector state, and we ensure stability in this state space. Given that the hammer is attached to the robot through a double hook, this assumption is valid as long as the hammer’s head is securely held. This assumption breaks down towards the end of the motion when the robot places the hammer on the table. However, since the final states of both the hammer and the robot are close, we observed that ensuring stability in the end-effector’s motion effectively guides the hammer’s motion towards its goal state, as can be seen in Fig. 12 and the attached video.
We used the Kinova Gen2 Ultra lightweight robot arm with a double hook replacing its default gripper. The control commands obtained from $f_{\theta}^{\mathcal{T}}$ were directly sent to a Cartesian space controller from Kinova [^8] Kinova’s controllers: https://github.com/Kinovarobotics/kinova-ros. The demonstrations were collected through kinesthetic teaching, i.e., physically moving the robot along desired paths while having the robot in gravity compensation mode. The simulated performance in this task was similar to that in the greenhouse, given that the state space was identical in both cases. We refer the reader to the attached video for examples of the robot executing this task from various initial conditions.
@@ 6 Conclusions
Lastly, while this paper’s findings are promising, there are also limitations that can be addressed in future works. Firstly, further exploration of the scalability properties of PUMA is needed, as the largest state space employed in this paper had 7 dimensions. Secondly, we have so far focused on learning independent motion primitives for specific tasks. A relevant line of research would be integrating this model into a larger framework where multiple primitives are learned and combined together. Finally, another interesting research direction is studying the integration of these approaches with the low-level control of the robots. Currently, it is assumed that the transitions requested by PUMA can be tracked by the controllers of the robot. However, this is not always the case, and it would be beneficial to incorporate this information into the learning framework.
@@ Acknowledgements
This research is funded by the Netherlands Organization for Scientific Research project Cognitive Robots for Flexible Agro-Food Technology, grant P17-01, and by the NXTGEN national program.
@@ Appendix A Upper Bound $\beta$
[…]
@@
[…]
Then, under the conditions of Theo. 4, i.e., $\forall t\in\mathbb{R}_{\geq 0}$,
@@
Figure 14: Summary of lemmas and corollaries related to the proof of Prop. 1. Bisque-colored boxes indicate the conditions necessary for the proposition to be true, while blue boxes represent lemmas relying directly on its assumptions and requirements.
@@
To demonstrate that this condition is valid, we introduce Lemma 1. For this lemma to be applicable, a specific condition must be met. To validate this condition, we refer to Lemma 2 and Cor. 1, both of which will be introduced subsequently. We also present Cor. 2, which will prove beneficial in subsequent discussions.
@@
    $d_{t}=d_{t+\Delta t}$,    for $y_{t+\Delta t}=y_{\mathrm{g}}$ (Theo. 4, 1),
@@
    $d_{t}>d_{t+\Delta t}$,    $\forall y_{t}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$ (Theo. 4, 2).
@@
[…]
@@
[…]
@@
Then, provided that the conditions of Prop. 1 are met, the conditions of Lemma 2 are fulfilled. Therefore, $\delta(y_{\mathrm{g}},t)=0$ for every $t\in\mathbb{R}_{\geq 0}$, and, hence, ${\delta^{\text{max}}(0,t)=0}$ (from Cor. 1). Consequently, we can employ Lemma 1 to demonstrate that $\beta(0,t)=0$ for all $t\in\mathbb{R}_{\geq 0}$.
@@
[…]
@@
[…]
@@
    $d_{t}>d_{t+\Delta t}$, $\forall t\in\mathbb{R}_{\geq 0}$, $\forall y_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$, (Theo. 4, 2).
@@
As a consequence, the assumptions and conditions outlined in Prop. 1 enable us to apply Lemma 4 to demonstrate that the conditions of Lemma 3 are satisfied. This, in turn, leads us to the conclusion that $\beta>\delta$ and $\beta(d_{0},t)$ strictly decreases with respect to $t$, $\forall d_{0}\in{\mathbb{R}_{\geq 0}\setminus\{0\}}$, $\forall t\in\mathbb{R}_{\geq 0}$. Lastly, before we showed that, for $d_{0}=0$, $\beta(0,t)=0$, $\forall t\in\mathbb{R}_{\geq 0}$. Therefore, in general, we have that $\beta\geq\delta$ and $\beta$ decreases with respect to $t$, $\forall d_{0}\in{\mathbb{R}_{\geq 0}}$, $\forall t\in\mathbb{R}_{\geq 0}$.
@@
[…]
@@
[…]
@@
    Continuity in $t$: For any given $d_{0}$, since $\beta$ evolves as a linear first-order system (as per (21)), its solution will exist provided that $\delta^{\text{max}}$ is continuous with respect to $t$ [55]. Then, under the assumptions of Prop. 1, Lemma 6 confirms that $\delta^{\text{max}}$ is indeed continuous, provided that $\delta$ is continuous. From Lemma 5 we infer that $\delta$ is continuous with respect to $t$. Hence, $\beta$ is well-defined for any given $d_{0}$, indicating that it is differentiable with respect to $t$, and, therefore, continuous.
@@
    Continuity in $d_{0}$: For any given $t$, from (21), we know that $\beta$ will be continuous with respect to $d_{0}$ as long as both of its terms, $z_{0}$ and the intregral of $\dot{z}_{t}$, are continuous. The continuity of both terms depends on the continuity of $\delta^{\text{max}}$. Then, since Lemma 7 indicates that $\delta^{\text{max}}$ is continuous with respect to $d_{0}$ provided that $\delta$ is continuous with respect to $y_{0}$, and Lemma 5 confirms that this is indeed the case (given the assumptions of Prop. 1), we conclude that $\beta$ is continuous with respect to $d_{0}$.
@@
To prove that this condition holds, we will utilize Lemma 8. The application of this lemma necessitates the use of Lemma 4 and Cor. 1, which are already introduced previously, and Lemma 9, which is introduced afterwards.
@@
[…]
@@
    $d_{t}<d_{t-\Delta t}$, $\forall y_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$, $\forall t\in\mathbb{R}_{\geq 0}$, (Theo. 4, 2).
@@
[…]
@@
To summarize, given the conditions and assumptions of Prop. 1, Lemmas 4, 9 and Cor. 1, indicate that the conditions of Lemma 8 are fulfilled, implying that $\beta\to 0$ as $t\to\infty$.
@@
From Lemma 10, we can conclude that $\beta$ strictly increases with respect to $d_{0}$, since assuming the conditions of Prop. 1, and that we have proven that $\beta$ is continuous in $t$, Lemmas 11, 3, and 8 indicate that the conditions of Lemma 10 are fulfilled.
To summarize, we have demonstrated that all the requirements for $\beta(d_{0},t)$ to be a class-$\mathcal{KL}$ function, and to serve as an upper bound for $\delta(y_{0},t)$, are met for every $y_{t}\in\mathcal{L}$ and for all $t\in\mathbb{R}_{\geq 0}$. With this, our proof of Prop. 1 is complete. ∎

### v3

@@ 5.2.1 Accuracy
Fig. 10a presents the performance of two variations of PUMA, namely when using Euclidean and spherical distance functions. Moreover, BC is included, serving as a lower-bound reference. As explained in Sec. 5.1.1, the boundary loss $\ell_{\partial}$ does not apply in this case and is, therefore, not evaluated. We can observe that the accuracy performance of both variations of PUMA is very similar, and BC performs slightly better than both of them.
@@ 5.2.2 Stability
Table 3 depicts the stability results for the LASA $\mathcal{S}^{2}$ dataset. In this case, we also conducted a stability analysis of CONDOR, which, as discussed in Sec. 2, should not be capable of ensuring stability in non-Euclidean state spaces. This assertion is confirmed by the data in Table 3, which shows that CONDOR yields 7.3133% of unsuccessful trajectories. When compared to BC, which has a 15.9217% rate of unsuccessful trajectories, it can be inferred that CONDOR reduces the number of unsuccessful trajectories in non-Euclidean state spaces. However, its performance is still far from satisfactory. In contrast, both versions of PUMA achieve 0% of unsuccessful trajectories, marking a significant improvement. Moreover, this validates that the Euclidean distance is effective for enforcing stability in spheres, as it can induce spherical metrics, such as the chordal distance, in lower-dimensional manifolds (see Appendix B).
@@ 5.2.3 Qualitative Analysis
Fig. 11 illustrates motions learned with PUMA in $\mathcal{S}^{2}$ using both spherical and Euclidean distances. Within the observed region of the sphere, the vector field exhibits no spurious attractors. Furthermore, when initiated from the same conditions as the demonstrations, the trajectories in both cases show an accurate reproduction of the motions.
@@ 5.2.4 State-of-the-art Comparison
Fig. 10b depicts a comparison between three methods: 1) CONDOR, 2) PUMA, and 3) Lie Flows [11], a method developed for learning stable motions in non-Euclidean manifolds using tools from Lie Theory. PUMA outperforms both CONDOR and Lie Flows in terms of accuracy. Moreover, in contrast to CONDOR, PUMA also successfully ensured stability in this dataset.
@@ 5.3 Real-world Experiments
We validate our method in two real-world setups, using two different robots. Both robots have six degrees of freedom, with their end-effector pose represented in $\mathbb{R}^{3}\times\mathcal{S}^{3}$ and are guided using PUMA. In these experiments, $f_{\theta}^{\mathcal{T}}$ is employed to generate velocity references in real-time, which are then sent to the low-level controllers responsible for tracking these references in the robots. Note that throughout the experiments, we operated under the assumption that the target states of the robots were already known.
@@ 5.3.1 Greenhouse
This experiment was conducted in a greenhouse, where a robot was trained to reach a black marker on a tomato plant (see Fig. 12). This marker, simulating a plant’s peduncle, represented the target state the robot must reach to harvest the plant. Here, the marker allowed us to test the method without causing harm to the plant. This task presents significant challenges, as it requires the robot to perform precise position and orientation control while considering the plant’s complex geometry, which is non-trivial to model. In a practical setting, a motion library should be developed to account for the variability across different plants and targets. This library could then be used to select or combine suitable motions to create an appropriate movement strategy considering the plant’s characteristics. This experiment represents a preliminary step towards achieving this objective.
We utilized an ABB IRB 1200 robot equipped with the Externally Guided Motion [^7] Code: https://github.com/ros-industrial/abb_robot_driver. module, allowing real-time joint velocity control. As the motion takes place in the end-effector space, the desired velocity at each time step was calculated using $f_{\theta}^{\mathcal{T}}$, and subsequently mapped to the joint space. This conversion was accomplished using the inverse kinematics module with joint limits from the Robotics Toolbox [52]. Similarly, we translated the robot’s estimated joint state into the end-effector space using the forward kinematics module from the same toolbox. The same approach was taken to collect the demonstrations of this task directly in the end-effector space, where a space mouse was employed to teleoperate the robot.
Fig. 13 presents the demonstrations and simulations of motions conducted in this experiment. It can be observed that five demonstrations were provided. Note that these demonstrations do not have the same end time; only the final state needs to be consistent. Over a span of 70 seconds, all simulated trajectories visibly converge to the goal. For more details, the reader is referred to the attached video.
@@ 5.3.2 Hammer
The second experiment involves a robot accurately positioning a hammer next to other tools on a table. This task requires precise control of the robot’s position and orientation, as shown in Fig. 12. We are interested in controlling the hammer’s movement, so we assume that the state of the hammer directly relates to the state of the robot’s end effector. This assumption is necessary because the employed low-level controller acts on the robot’s end-effector state, and we ensure stability in this state space. Given that the hammer is attached to the robot through a double hook, this assumption is valid as long as the hammer’s head is securely held. This assumption breaks down towards the end of the motion when the robot places the hammer on the table. However, since the final states of both the hammer and the robot are similar, we observed that ensuring stability in the end-effector’s motion effectively guides the hammer’s motion towards its goal state, as can be seen in Fig. 12 and in the attached video.
We used the Kinova Gen2 Ultra lightweight robot arm with a double hook replacing its default gripper. The control commands obtained from $f_{\theta}^{\mathcal{T}}$ were directly sent to a Cartesian space controller from Kinova [^8] Kinova’s controllers: https://github.com/Kinovarobotics/kinova-ros. The demonstrations were collected through kinesthetic teaching, i.e., physically moving the robot along desired paths. The simulated performance in this task was similar to that in the greenhouse, given that the state space was identical in both cases. We refer the reader to the attached video for examples of the robot executing this task from various initial conditions.
@@ 6 Conclusions
Lastly, while this paper’s findings are promising, there are also limitations that can be addressed in future works. First, further exploration of the scalability properties of PUMA is needed, as the largest DNN input employed in this paper had 7 dimensions. Second, we have so far focused on learning independent motion primitives for specific tasks. A relevant line of research would be integrating this model into a larger framework where multiple primitives are learned and combined together. Third, a topic that we did not address in this work is obstacle avoidance. Recent techniques, such as Geometric Fabrics [53], exploit dynamical systems to represent motions that achieve full-body obstacle avoidance, which can be integrated with dynamical systems learned in the end-effector space of the robot. Consequently, an exciting avenue for future work is exploring the combination of these methods with PUMA. Finally, another interesting research direction is studying the integration of these approaches with the low-level control of the robots. Currently, it is assumed that the transitions requested by PUMA can be tracked by the controllers of the robot. However, this is not always the case, and it would be beneficial to incorporate this information into the learning framework.
@@ Acknowledgements
This research is funded by the Netherlands Organization for Scientific Research project Cognitive Robots for Flexible Agro-Food Technology, grant P17-01, and by the National Growth Fund program NXTGEN Hightech.
@@ Appendix A Upper Bound $\beta$
$\delta^{\text{max}}(d_{0},t)=\max_{y_{0}\in\mathcal{Y}_{0}}\left(\max_{s\in[t,t+\Delta t]}\delta(y_{0},s)\right),$ (23)
@@
$z_{0}=\delta^{\text{max}}_{0}+d_{0},$ (24)
where $\delta^{\text{max}}_{0}=\delta^{\text{max}}(d_{0},0)$.
Then, under the conditions of Theorem 4, i.e., $\forall t\in\mathbb{R}_{\geq 0}$,
@@
Figure 14: Summary of lemmas and corollaries related to the proof of Proposition 1. Bisque-colored boxes indicate the conditions necessary for the proposition to be true, while blue boxes represent lemmas relying directly on its assumptions and requirements.
@@
To demonstrate that this condition is valid, we introduce Lemma 1. For this lemma to be applicable, a specific condition must be met. To validate this condition, we refer to Lemma 2 and Corollary 1, both of which will be introduced subsequently. We also present Corollary 2, which will prove beneficial in subsequent discussions.
@@
    $d_{t}=d_{t+\Delta t}$,    for $y_{t+\Delta t}=y_{\mathrm{g}}$ (Theorem 4, 1),
@@
    $d_{t}>d_{t+\Delta t}$,    $\forall y_{t}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$ (Theorem 4, 2).
@@
${\delta^{\text{max}}(0,t)=0}$ if ${\delta(y_{\mathrm{g}},t)=0}$, $\forall t\in\mathbb{R}_{\geq 0}$.
@@
$\min(\delta^{\text{max}})=0$ if ${\delta^{\text{max}}(0,t)=0}$, $\forall t\in\mathbb{R}_{\geq 0}$.
@@
Then, provided that the conditions of Proposition 1 are met, the conditions of Lemma 2 are fulfilled. Therefore, $\delta(y_{\mathrm{g}},t)=0$ for every $t\in\mathbb{R}_{\geq 0}$, and, hence, ${\delta^{\text{max}}(0,t)=0}$ (from Corollary 1). Consequently, we can employ Lemma 1 to demonstrate that $\beta(0,t)=0$ for all $t\in\mathbb{R}_{\geq 0}$.
@@
    $\delta^{\text{max}}$ decreases with respect to $t$, $\forall d_{0}\in\mathbb{R}_{\geq 0}\setminus\{0\}$.
@@
$\delta^{\text{max}}(d_{0},t)$ decreases with respect to $t$, $\forall d_{0}\in\mathbb{R}_{\geq 0}\setminus\{0\}$, $\forall t\in\mathbb{R}_{\geq 0}$, if:
@@
    $d_{t}>d_{t+\Delta t}$, $\forall t\in\mathbb{R}_{\geq 0}$, $\forall y_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$, (Theorem 4, 2).
@@
As a consequence, the assumptions and conditions outlined in Proposition 1 enable us to apply Lemma 4 to demonstrate that the conditions of Lemma 3 are satisfied. This, in turn, leads us to the conclusion that $\beta>\delta$ and $\beta(d_{0},t)$ strictly decreases with respect to $t$, $\forall d_{0}\in{\mathbb{R}_{\geq 0}\setminus\{0\}}$, $\forall t\in\mathbb{R}_{\geq 0}$. Lastly, before we showed that, for $d_{0}=0$, $\beta(0,t)=0$, $\forall t\in\mathbb{R}_{\geq 0}$. Therefore, in general, we have that $\beta\geq\delta$ and $\beta$ decreases with respect to $t$, $\forall d_{0}\in{\mathbb{R}_{\geq 0}}$, $\forall t\in\mathbb{R}_{\geq 0}$.
@@
$\delta^{\text{max}}$ is a continuous function with respect to $t$, if $\delta$ is continuous with respect to $t$.
@@
$\delta^{\text{max}}$ is continuous with respect to $d_{0}$, if $\delta$ is continuous with respect to $y_{0}$.
@@
    Continuity in $t$: For any given $d_{0}$, since $\beta$ evolves as a linear first-order system (as per (21)), its solution will exist provided that $\delta^{\text{max}}$ is continuous with respect to $t$ [58]. Then, under the assumptions of Proposition 1, Lemma 6 confirms that $\delta^{\text{max}}$ is indeed continuous, provided that $\delta$ is continuous. From Lemma 5 we infer that $\delta$ is continuous with respect to $t$. Hence, $\beta$ is well-defined for any given $d_{0}$, indicating that it is differentiable with respect to $t$, and, therefore, continuous.
@@
    Continuity in $d_{0}$: For any given $t$, from (21), we know that $\beta$ will be continuous with respect to $d_{0}$ as long as both of its terms, $z_{0}$ and the intregral of $\dot{z}_{t}$, are continuous. The continuity of both terms depends on the continuity of $\delta^{\text{max}}$. Then, since Lemma 7 indicates that $\delta^{\text{max}}$ is continuous with respect to $d_{0}$ provided that $\delta$ is continuous with respect to $y_{0}$, and Lemma 5 confirms that this is indeed the case (given the assumptions of Proposition 1), we conclude that $\beta$ is continuous with respect to $d_{0}$.
@@
To prove that this condition holds, we will utilize Lemma 8. The application of this lemma necessitates the use of Lemma 4 and Corollary 1, which are already introduced previously, and Lemma 9, which is introduced afterwards.
@@
    $\delta^{\text{max}}$ decreases with respect to $t$, $\forall d_{0}\in\mathbb{R}_{\geq 0}$, $\forall t\in\mathbb{R}_{\geq 0}$,
@@
    $d_{t}<d_{t-\Delta t}$, $\forall y_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$, $\forall t\in\mathbb{R}_{\geq 0}$, (Theorem 4, 2).
@@
$\delta^{\text{max}}\in[0,\max(\delta)]$, and $\delta^{\text{max}}$ surjective with respect to $d_{0}$, if:
@@
To summarize, given the conditions and assumptions of Proposition 1, Lemmas 4, 9 and Corollary 1, indicate that the conditions of Lemma 8 are fulfilled, implying that $\beta\to 0$ as $t\to\infty$.
@@
From Lemma 10, we can conclude that $\beta$ strictly increases with respect to $d_{0}$, since assuming the conditions of Proposition 1, and that we have proven that $\beta$ is continuous in $t$, Lemmas 11, 3, and 8 indicate that the conditions of Lemma 10 are fulfilled.
To summarize, we have demonstrated that all the requirements for $\beta(d_{0},t)$ to be a class-$\mathcal{KL}$ function, and to serve as an upper bound for $\delta(y_{0},t)$, are met for every $y_{t}\in\mathcal{L}$ and for all $t\in\mathbb{R}_{\geq 0}$. With this, our proof of Proposition 1 is complete. ∎